OpenStudy (anonymous):

Given the function y = 5x – 7, how does the y-value change as x increases by 1? A. The y-value decreases by 7. B. The y-value decreases by 2. C. The y-value increases by 12. D. The y-value increases by 5.

OpenStudy (ivettef365):

plug the number on x to see how much increases or decreases as follows: you have now y = 5x -7 that means (0,7) y = 5(1)- 7 = 5-7 = -2 (1,-2) y = 5(2) -7 = 10 -7 (2,3)

OpenStudy (ivettef365):

so what is happening to y when x increases by 1

OpenStudy (ivettef365):

I meant the first one should be (0,-7)

OpenStudy (anonymous):

im confuse at the max

OpenStudy (ivettef365):

the range (y) {-7,-2,3......} what is the sequence

OpenStudy (anonymous):

+5 right?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

yay!!!!

OpenStudy (anonymous):

so D